# Transfer-Based Strategies for Multi-Target Quantum Optimization

### v1

@@ V-B Limitation based on the number of qubits
where the distance between any two target $\{x,y\}$ is simply defined as the norm-2 $\|y-x\|$. Figure 4 illustrates the expected minimum distance between a new target $y\sim\mathcal{B}$ and a set of $K$ previously targets $x\in\mathcal{S}$. The vertical axis shows the quantity $\mathbb{E}_{y\sim B}\left[\min_{x\in A}\|y-x\|\right]$, measuring how close a new target is, on average, to its nearest neighbor in $A$. The horizontal axis corresponds to $K=|A|$. Across all cases from 2 to 8-qubit, increasing $K$ leads to a monotonic decrease in the expected minimum distance, as a larger reference set increases the probability of locating a closer neighbor. However, this decay saturates as $K$ grows, indicating diminishing returns with additional targets.
Furthermore, the distances increase with the number of qubits. For example, the 2-qubit case achieves significantly lower average distances than the 8-qubit case at all values of $k$. This trend is consistent with the curse of dimensionality: as the dimension of $\mathcal{S}$ increases, random vectors become increasingly distant, leading to sparser convergence of the space even at higher $K$. The error bars are largest at low $K$, especially for lower-qubit configurations, reflecting high variability when the reference set is sparse. These results underscore the growing challenge of generalization and interpolation in high-dimensional quantum state spaces. In particular, nearest-neighbor-based retrieval becomes less effective with increasing qubit number, motivating the need for more expressive models. So on, the warm-starting method will not be suitable for large-qubit optimization because $K$ must be increased exponentially based on the dimension of $\mathcal{S}$.
@@ VI Conclusion
Our results suggest that the effectiveness of transfer increases with the number of pre-optimized targets, and that using appropriate similarity metrics is critical for effective knowledge transfer. However, the diminishing returns observed in higher-dimensional cases, due to the curse of dimensionality, indicate limitations when scaling to larger systems. Consequently, while warm-start and neighbor-based estimators are practical for small- and mid-sized circuits, alternative strategies such as meta-learning or hierarchical learning are needed for high-qubit regimes.
@@
*Hidden for peer review

### v2

@@ V-B Limitation based on the number of qubits
where the distance between any two target $\{x,y\}$ is simply defined as the norm-2, $\|y-x\|\equiv\|y-x\|_{2}=2-2\mathbf{Re}(\langle x|y\rangle)$. The distance metric is norm-2 instead of fidelity in case we want to control the density of the generated dataset. Figure 4 illustrates the expected minimum distance between a new target $y\sim\mathcal{B}$ and a set of $K$ previously targets $x\in\mathcal{S}$. The vertical axis shows the quantity $\mathbb{E}_{y\sim B}\left[\min_{x\in A}\|y-x\|\right]$, measuring how close a new target is, on average, to its nearest neighbor in $A$. The horizontal axis corresponds to $K=|A|$. Across all cases from 2 to 8-qubit, increasing $K$ leads to a monotonic decrease in the expected minimum distance, as a larger reference set increases the probability of locating a closer neighbor. However, this decay saturates as $K$ grows, indicating diminishing returns with additional targets.
Furthermore, the distances increase with the number of qubits. For example, the 2-qubit case achieves significantly lower average distances than the 8-qubit case at all values of $k$. This trend is consistent with the curse of dimensionality: as the dimension of $\mathcal{S}$ increases, random vectors become increasingly distant, leading to sparser convergence of the space even at higher $K$. The error bars are largest at low $K$, especially for lower-qubit configurations, reflecting high variability when the reference set is sparse. These results underscore the growing challenge of generalization and interpolation in high-dimensional quantum state spaces. In particular, nearest-neighbor-based retrieval becomes less effective with increasing qubit number, motivating the need for more expressive models. So, the warm-starting method becomes unsuitable for large-qubit optimization since $K$ must increase exponentially with the dimension of $S$.
@@ VI Conclusion
Our results suggest that the effectiveness of transfer increases with the number of pre-optimized targets and that using appropriate similarity metrics is critical for effective knowledge transfer. However, the diminishing returns observed in higher-dimensional cases, due to the curse of dimensionality, indicate limitations when scaling to larger systems. Consequently, while warm-start and neighbor-based estimators are practical for small- and mid-sized circuits, alternative strategies such as meta-learning or hierarchical learning are needed for high-qubit regimes.
@@
The codes used for this study are available at https://github.com/NAIST-Archlab/MTQO.